Johnson Products is considering purchasing a new milling machine that costs $120,000. The machine’s installation and shipping costs will total $4,500. If accepted, the milling machine project will require an initial net working capital investment of $20,000. Johnson plans to depreciate the machine on a straight-line basis over a period of 8 years. About a year ago, Johnson paid $12,000 to a consulting firm to conduct a feasibility study of the new milling machine. Johnson’s marginal tax rate is 40 percent.

  1. Calculate the project’s net investment (NINV). Round your answer to the nearest dollar.

    $  

  2. Calculate the annual straight-line depreciation for the project. Round your answer to the nearest cent.

    $  

  3. Calculate MACRS depreciation assuming this is a 7-year class asset. Use Table 9A-3 to answer the question. Round your answers to the nearest dollar.

    Year Depreciation
    1 $  
    2 $  
    3 $  
    4 $  
    5 $  
    6 $  
    7 $  
    8 $  
Answer #1

a) New machine cost Add: Installation & shipping cost Add: Working capital $ $ $ 120,000 4,500 20,000 Net investment $ 144,50

2 All 3 Year Depreciation (MACRS Rate*Asset cost) 1 $ 17,791.05 =124500*14.29% $ 30,490.05 =124500*24.49% $ 21,775.05 =124500

*Please rate thumbs up